# A Heuristic Algorithm Based on Beam Search and Iterated Local Search for the Maritime Inventory Routing Problem

### v1

@@ 5.3 Results Analysis
This section presents the results of the tests conducted to assess the performance of the proposed methodologies.
@@ 5.3.1 BS and Greedy Heuristic Results
Originally, this greedy heuristic served as the sole constructive approach but was unable to produce high-quality solutions, as can be observed in Table 8 (7).
As mentioned in Section 4.3, various combinations of nondeterminism were tested, applying it separately to port selection, vessel selection, or both. Thus, by incorporating a probability factor within a normal distribution and considering standard deviation, it is possible that the next port on the route might not be the first to face risk, and the vessel selected may not be the one that reaches there first. The normal distribution was used to sample values around the average, introducing variation while keeping most variability but ensuring that the majority of samples close to that remained near the central value. The configuration that achieved the best result was to apply randomness only to port selection, while maintaining deterministic vessel selection. For these tests, the 13 smallest Group 2 instances were used, with 4 different seeds, all with a 120-day horizon, and $q$ was set to 5. The size of these instances made it possible to set a large value for $N$, namely $N=1000$. This balances high-quality results with acceptable CPU time. Table 9 (7) presents the results of these tests.
Once it was established that nondeterministic methods would be used solely for the port, additional tests were performed to evaluate three distinct techniques for aggregating the results of the greedy algorithm: namely, maximum value, mean, and median. Experiments were then conducted for the 13 smallest instances using three horizon sizes. For each instance, four executions with different seeds were carried out. The number of nodes allowed per level in the BS was set to 200. This decrease is attributed to the fact that instances now take into account a time frame extending up to 360 days. The method that showed the best performance was the median. The results of these tests are shown in Table 10 (7).
Table 3 highlights the significance of the BS within the algorithm. Each column corresponds to a specific algorithm stage (After BS, After LS, and After ILS), allowing a comparison of the impact of BS on the results. The percentages indicate the improvement achieved by incorporating BS as opposed to relying solely on the greedy heuristic. Comparing the results obtained through the exclusive use of greedy heuristic versus the application of the BS algorithm, it is possible to observe that the BS approach yields an average improvement of 18.40% in the solution following the BS stage. Notably, local search has a significant impact on the results in the absence of BS, reducing the advantage of employing BS to 7.14%. When the complete heuristic is applied, the benefit of BS further decreases to 3.57% for this set of instances. However, its use remains advantageous, despite the increased computational effort.
@@
5.3.2 Proposed Heuristic Results
Two studies (Papageorgiou, 2013; Papageorgiou et al., 2018) made use of MIRPLib Group 2 instances. In the first study (Papageorgiou, 2013), the authors first performed a warm-up phase to obtain an initial good solution, which was then used to warm-start the solver. Subsequently, the main experiments were run for 24 hours. In the subsequent investigation (Papageorgiou et al., 2018), the overall CPU time per scenario remains ambiguous, due to its combination of various techniques with different time periods. Papageorgiou (2013), however, emphasizes that it is difficult to find good solutions for this group in a few minutes. There is a balance between the quality of the results and the time required to obtain them. Therefore, this work aims to evaluate how heuristic works under different CPU time conditions. To accomplish this, the parameter $N$ for the BS is evaluated using three distinct values ($N$ = 10, $N$ = 100, $N$ = 1000). For each $N$, if the instance does not reach 24 hours of CPU time, it moves to the next value of $N$. If it exceeds 24 hours, it will not proceed to the next $N$ value. This approach is intended to analyze how results improve with increased CPU time using the work by Papageorgiou et al. (2018) as a baseline reference. Another important point is that the gap reported in the results below is the best gap obtained from the runs executed within 90,000 seconds. If any stage of the main algorithm was not completed within this time frame, the reported result corresponds to the last completed stage. For example, if the algorithm does not finish the ILS phase, the reported value reflects the best result achieved up to the LS, along with its respective time. Similarly, if the algorithm does not complete the LS, the reported result comes from the BS, including both the objective value and the time.
The main results for the 120, 180, and 360 periods are summarized in Tables 5, 6, and 7, respectively. The tables include key metrics such as the value of $N$, the classification of each instance, the objective values (Objval), the total costs, the percentage of the objective gap (GAP Objval) and the CPU time (in hours). These tables show the best results achieved in 90,000 seconds (approximately 25 hours). The reported results represent the best performances obtained across all the values tested for $N$. The cases indicated by (*) show results for BS and LS, as the ILS did not complete within 90,000 seconds. The proposed heuristic demonstrated improvements in the objective value of MIRPLib (Papageorgiou et al., 2018) in ten different instances when $N$ = 1000. As expected, increasing the number of nodes improves the results. However, this improvement is accompanied by a trade-off, specifically an increase in CPU time. Furthermore, regarding the classification of instances, it can be observed that those classified as “E” generally exhibit lower objective gaps compared to instances classified as “M” or “H”.
Table 5: 120-period results
Instance N Classification Objval Total Cost GAP Objval (%) Time (h)
LR1_DR02_VC01_V6a 1000 E 33809.00 33808.95 0.00 0.03
LR1_DR02_VC02_V6a 1000 E 74982.00 74986.52 0.01 0.04
LR1_DR02_VC03_V7a 1000 E 40446.00 40403.90 $-0.10$ 0.05
LR1_DR02_VC03_V8a 1000 E 43721.00 43933.34 0.48 0.04
LR1_DR02_VC04_V8a 1000 E 41657.00 41847.57 0.46 0.08
LR1_DR02_VC05_V8a 1000 E 36659.00 36616.09 $-0.12$ 0.06
LR1_DR03_VC03_V10b 1000 M 92941.00 95667.58 2.85 0.18
LR1_DR03_VC03_V13b 1000 E 124921.00 127203.29 1.79 0.43
LR1_DR03_VC03_V16a 1000 E 82837.00 91335.26 9.30 0.26
LR1_DR04_VC03_V15a 1000 E 73312.00 74714.08 1.88 0.67
LR1_DR04_VC03_V15b 1000 E 117812.00 116451.06 $-1.17$ 0.77
LR1_DR04_VC05_V17a 1000 E 72876.00 73864.03 1.34 0.89
LR1_DR04_VC05_V17b 1000 E 105766.00 103534.03 $-2.16$ 1.02
LR1_DR05_VC05_V25a 1000 E 105328.00 105995.54 0.63 2.57
LR1_DR05_VC05_V25b 1000 E 137107.00 140504.31 2.42 2.87
LR1_DR08_VC05_V38a 1000 M 166615.00 169501.81 1.70 19.83
LR1_DR08_VC05_V40a* 1000 E 178593.00 185214.23 3.57 24.94
LR1_DR08_VC05_V40b 100 M 200746.00 234070.81 14.24 4.46
LR1_DR08_VC10_V40a 100 M 185538.00 189991.43 2.34 3.43
LR1_DR08_VC10_V40b 100 M 206315.00 234282.23 11.94 4.40
LR1_DR12_VC05_V70a* 100 H 278647.00 289736.67 3.83 18.37
LR1_DR12_VC05_V70b* 100 M 308555.00 322301.84 4.27 19.98
LR1_DR12_VC10_V70a 10 H 283154.00 290673.25 2.59 12.61
LR1_DR12_VC10_V70b 10 H 295126.00 304041.16 2.93 15.71
Table 6: 180-period results
Instance N Classification Objval Total Cost GAP Objval (%) Time (h)
LR1_DR02_VC01_V6a 1000 E 52167.00 52167.21 0.00 0.07
LR1_DR02_VC02_V6a 1000 E 129372.00 129696.36 0.25 0.09
LR1_DR02_VC03_V7a 1000 M 60547.00 61950.19 2.27 0.11
LR1_DR02_VC03_V8a 1000 E 68153.00 70478.74 3.30 0.09
LR1_DR02_VC04_V8a 1000 M 66017.00 66238.44 0.33 0.17
LR1_DR02_VC05_V8a 1000 M 58619.00 58202.65 $-0.72$ 0.14
LR1_DR03_VC03_V10b 1000 M 125638.00 122230.31 $-2.79$ 0.38
LR1_DR03_VC03_V13b 1000 E 165764.00 161724.70 $-2.50$ 0.90
LR1_DR03_VC03_V16a 1000 M 143178.00 153068.94 6.46 0.62
LR1_DR04_VC03_V15a 1000 M 118621.00 119146.56 0.44 1.50
LR1_DR04_VC03_V15b 1000 E 189989.00 188776.86 $-0.64$ 1.66
LR1_DR04_VC05_V17a 1000 M 117710.00 118868.81 0.97 2.10
LR1_DR04_VC05_V17b 1000 M 159168.00 154452.82 $-3.05$ 2.27
LR1_DR05_VC05_V25a 1000 M 171620.00 172104.96 0.28 6.04
LR1_DR05_VC05_V25b 1000 M 205368.00 219522.01 6.45 6.57
LR1_DR08_VC05_V38a 100 H 274244.00 281247.89 2.49 9.00
LR1_DR08_VC05_V40a 100 H 296760.00 304072.84 2.40 9.77
LR1_DR08_VC05_V40b 100 M 337559.00 354956.21 4.90 11.61
LR1_DR08_VC10_V40a 100 H 304261.00 309323.52 1.64 9.41
LR1_DR08_VC10_V40b 100 H 331775.00 356027.44 6.81 11.99
LR1_DR12_VC05_V70a* 10 H 460566.00 506963.62 9.15 3.40
LR1_DR12_VC05_V70b* 10 H 491160.00 573240.49 14.32 3.77
LR1_DR12_VC10_V70a* 10 H 466975.00 491159.69 4.92 3.77
LR1_DR12_VC10_V70b* 10 H 470172.00 511421.68 8.07 4.18
Table 7: 360-period results
Instance N Classification Objval Total Cost GAP Objval (%) Time (h)
LR1_DR02_VC01_V6a 1000 E 108141.00 108141.00 0.00 0.31
LR1_DR02_VC02_V6a 1000 E 283031.00 283788.36 0.27 0.44
LR1_DR02_VC03_V7a 1000 H 124282.00 126940.04 2.09 0.50
LR1_DR02_VC03_V8a 1000 M 141166.00 146807.48 3.84 0.41
LR1_DR02_VC04_V8a 1000 M 138693.00 139237.84 0.39 0.84
LR1_DR02_VC05_V8a 1000 H 122598.00 122859.71 0.21 0.71
LR1_DR03_VC03_V10b 1000 H 259888.00 260807.85 0.35 1.56
LR1_DR03_VC03_V13b 1000 N 316441.00 312621.28 $-1.22$ 3.88
LR1_DR03_VC03_V16a 1000 N 327793.00 349216.16 6.13 3.56
LR1_DR04_VC03_V15a 1000 N 252710.00 256288.64 1.40 7.75
LR1_DR04_VC03_V15b 1000 N 339308.00 340710.34 0.41 7.63
LR1_DR04_VC05_V17a 1000 H 251623.00 254040.64 0.95 11.23
LR1_DR04_VC05_V17b 1000 M 304507.00 312046.69 2.42 10.18
LR1_DR05_VC05_V25a 100 H 368628.00 373871.33 1.40 8.84
LR1_DR05_VC05_V25b 100 H 410053.00 494627.59 17.10 9.79
LR1_DR08_VC05_V38a* 100 H 596969.00 683046.79 12.60 22.52
LR1_DR08_VC05_V40a* 100 H 652380.00 659967.40 1.15 20.30
LR1_DR08_VC05_V40b* 100 H 709713.00 804362.42 11.77 22.41
LR1_DR08_VC10_V40a* 100 H 663245.00 672938.28 1.44 19.92
LR1_DR08_VC10_V40b* 100 H 724513.00 794420.68 8.80 21.26
LR1_DR12_VC05_V70a* 10 H 1021389.00 1193627.62 14.43 20.07
LR1_DR12_VC05_V70b* 10 H 1093013.00 1435397.61 23.85 21.35
LR1_DR12_VC10_V70a* 10 H 1024399.00 1046678.91 2.13 19.39
LR1_DR12_VC10_V70b* 10 H 1001541.00 1306455.89 23.34 16.12
The box plots presented in Figures 5, 6, and 7 illustrate the relationship between different stages (BS, LS, and ILS) and the variable $N$ over three distinct values (10, 100, and 1000), providing relevant information on the influence of the parameter $N$ on the results. The y-axis represents the percentage of the objective gap (GAP Objval), while the x-axis denotes the varying values of $N$.
A general observation indicates that as the value of $N$ increases from 10 to 1000, there is a visible decrease in the median GAP Objval across all stages, particularly in the BS stage, which is directly affected by this parameter. This suggests that larger values of $N$ are associated with lower percentages of gaps. Furthermore, it can be concluded that while LS has a considerable influence in all the cases analyzed, whereas the effect of ILS tends to decrease as the value of $N$ increases. This observation remains consistent across various instance sizes, with the only notable variation being the increase in GAP Objval for larger instances. Finally, Additionally, the interquartile ranges for each stage illustrate that there is a significant reduction in variability as $N$ increases. At $N$ = 10, all stages exhibit a wide spread in values, with ILS showing the least variability. As $N$ increases, the number of outliers decreases, which indicates more consistent performance across stages.
Figure 5: H = 120 with different $N$ values
Figure 6: H = 180 with different $N$ values
Figure 7: H = 360 with different $N$ values
The box plots presented in Figures 8 and 9 illustrate the distribution of GapObjval and Time (h) across different values of $N$ (10, 100, 1000) and instance classifications (E, M, H). For GapObjval, instances classified as “E” consistently exhibit lower values, the median decreasing as $N$ increases. These instances are relatively easier to solve, achieving near-optimal solutions with a lower deviation. The class “M” presents higher variability, especially for smaller values of $N$, with a gradual reduction in dispersion as the size of the problem increases. In contrast, class “H” maintains a moderate but stable GapObjval distribution, with values slightly decreasing as $N$ grows, although with less visible variability compared to class “M”.
Figure 8: GAPObjval by Class and N values
Figure 9: Time by Class and N values
With respect to time (h), there is a clear increase in computational effort as $N$ grows. For Class “E”, CPU times remain low across all values of $N$, indicating efficient convergence. In Class “M”, CPU times increase significantly, especially for larger $N$ values ($N$ = 1000), where a wider spread suggests varying solver performance across different instances. The class “H” consistently requires the highest computational effort, and as $N$ grows, there is a substantial increase in CPU time, ratifying the difficulty of these instances. The CPU time distribution in Class “H” also exhibits larger variability, suggesting that some instances require substantially more computational resources.
As one of the challenges of MIRPLib Group 2 instances was achieving good results in a short time, we evaluated the GAPObjval over time to determine the results achieved after 3600 seconds, as Munguía et al. (2019) did in their work. Figures 10 to 13 show the results for easy, medium, hard and all instances.
In a first analysis, the results may indicate that Munguía’s method generally delivers better performance than ours. Below, it is presented the average performance across different instance categories and execution times (500 and 3600 seconds). Munguía’s results can be found in Munguía et al. (2019).
For easy instances, Figure 10, the gaps are smaller. At 500 seconds, Munguía et al. (2019) reached 3%, while our method had 5%. After 3600 seconds, the gaps decreased further to 2% (Munguía et al., 2019) and 4% (ours), maintaining the same relative difference.
Figure 10: GAPObjval over time and N values - Easy instances
For medium instances, Figure 11, the gaps were 10% (Munguía) and 15% (ours) at 500 seconds. At 3600 seconds, the values dropped to 5% and 11%, respectively, with Munguía et al. (2019) still showing better performance.
Figure 11: GAPObjval over time and N values - Medium instances
As expected, the hard instances yielded the highest gaps. At 500 seconds, Munguía et al. (2019) had 25%, and our method 27%. With extended execution (3600 seconds), both improved: 10% for Munguía et al. (2019) and 22% for ours.
Figure 12: GAPObjval over time and N values - Hard instances
Considering all instances, at 500 seconds, Munguía et al. (2019) achieved an average gap of 13%, while our method performed with 17%. After 3600 seconds, both methods improved significantly, but Munguía et al. (2019) maintained the advantage with 5%, compared to 13% for our method.
Figure 13: GAPObjval over time and N values - All instances
However, the observed performance gap is influenced by the computational environments in which each method was executed. While Munguía et al. (2019) ran their experiments on a 96-core Xeon X5650 cluster, our method was executed on a single thread of an i7-8700K processor. Each core of the i7-8700K delivers approximately 2.5 times the performance of a Xeon X5650 core (PassMark Software, 2024). Based on this, the total computational power available to Munguía’s method was 38 times greater than ours. Therefore, what takes us one hour to execute could be completed on their system in about 94 seconds. This means that the performance gap (GAPObjval) must be interpreted with caution. Probably, under comparable hardware conditions, our method could achieve competitive results.

### v2

@@ 5.3 Results Analysis
5.3.1 Proposed Heuristic Results
Two studies (Papageorgiou, 2013; Papageorgiou et al., 2018) used MIRPLib Group 2 instances. In the first study, the authors first performed a warm-up phase to obtain an initial good solution, which was then used to warm-start the solver. Subsequently, the main experiments were carried out for 24 hours. In the following research (Papageorgiou et al., 2018), the overall CPU time per scenario remains unclear, due to the combination of multiple techniques with different times. Moreover, Papageorgiou (2013) emphasizes that it is difficult to find good solutions for Group 2 in a few minutes, establishing a balance between the quality of the results and the time required to obtain them. Therefore, this work aims to evaluate how the heuristic works under different CPU times. To accomplish this, the parameter $N$ for the BS is evaluated using three different values ($N\in\{10,100,1000\}$). For each $N$, if the instance does not reach 24 hours of CPU time, it moves to the next value of $N$. If it exceeds 24 hours, it does not proceed to the next value of $N$. This approach is intended to analyze how results improve with increased CPU time using the work by Papageorgiou et al. (2018) as a baseline reference. Another important point is that the gap reported in our results is the best gap obtained from the executed runs in 90,000 seconds. If any stage of the main algorithm was not completed within this time limit, the reported result corresponds to the last completed stage. For example, if the algorithm does not finish the ILS phase, the reported value reflects the best result achieved up to the LS, along with its respective time. Similarly, if the algorithm does not complete the LS, the reported result comes from the BS, including both the objective value and the time.
The box plots presented in Figures 3, 4, and 5 illustrate the comparison between different stages (BS, LS, and ILS) and the variable $N$ over three distinct values (10, 100, and 1000), providing relevant information on the influence of the parameter $N$ on the results. The y-axis represents the percentage of the objective gap (Gap Obj), while the x-axis denotes the different values of $N$.
A general observation indicates that as the value of $N$ increases from 10 to 1000, there is a visible decrease in the median Gap Obj across all stages, particularly in the BS stage, which is directly affected by this parameter. This suggests that larger values of $N$ are associated with lower percentages of gaps. Furthermore, it can be concluded that while LS has a considerable influence in all the cases analyzed, the effect of ILS tends to decrease as the value of $N$ increases. This observation remains consistent across various instance sizes, with the only notable variation being the increase in Gap Obj for larger instances. Furthermore, the interquartile ranges for each stage illustrate that there is a significant reduction in variability as $N$ increases. At $N=10$, all stages exhibit a wide spread in values, with ILS showing the smallest variability. As $N$ increases, the number of outliers decreases, which indicates more consistent performance between stages.
Figure 3: H = 120 with different $N$.
Figure 4: H = 180 with different $N$.
Figure 5: H = 360 with different $N$.
The box plots presented in Figures 6 and 7 illustrate the distribution of Gap Obj and time between different values of $N$ and the classifications of instances. For Gap Obj, instances classified as easy consistently exhibit lower values, the median decreasing as $N$ increases. These instances are relatively easier to solve, achieving near-optimal solutions with a lower deviation. The medium class presents higher variability, especially for lower values of $N$, with a gradual reduction in dispersion as the size of the problem increases. In contrast, the hard class maintains a moderate but stable Gap Obj distribution, with values slightly decreasing as $N$ grows, although with less visible variability compared to the medium class.
Figure 6: Gap Obj by class and $N$.
With respect to time, there is a clear increase in computational effort as $N$ grows. For the easy class, CPU times remain low in all values of $N$, indicating efficient convergence. In the medium class, CPU times increase significantly, especially for higher $N$ values ($N$ = 1000), where a wider spread suggests inconsistent solver performance between different instances. The hard class consistently requires the highest computational effort, and as $N$ grows, there is a substantial increase in CPU time, confirming the difficulty of these instances. The CPU time distribution in the hard class also exhibits larger variability, suggesting that some instances require substantially more computational resources.
Figure 7: Time by Class and N.
The main results for instances with 120, 180, and 360 periods are summarized in Tables 3, 4, and 5, respectively. The tables include key metrics such as the value of $N$, the classification of each instance (Class), the objective values (Obj), the best total cost among the ten seeds (Best Total Cost), the best percentage of the objective gap (Best Gap), the average total costs (Average Total Cost), the average percentage of the objective gap (Average Gap) and the CPU time (in hours). These tables show the best results obtained in 90,000 seconds (approximately 25 hours). The reported results represent the best values obtained across all values tested for $N$. The cases indicated by (*) show results for BS and LS, as the ILS did not complete within 90,000 seconds. The proposed heuristic demonstrated improvements in the objective value of MIRPLib (Papageorgiou et al., 2018) in 19 different instances when $N$ = 1000, considering the best seed of each instance. As expected, increasing the number of nodes improves the results. However, this improvement is accompanied by a trade-off, specifically an increase in CPU time. Furthermore, regarding the classification of instances, it can be observed that those classified as easy generally exhibit lower objective gaps compared to those classified as medium or hard. Complete results can be found in the appendix.
Table 3: 120-period results
Best Best Average Average Time
Instance N Class Obj Total Cost Gap (%) Total Cost Gap (%) (h)
LR1_DR02_VC01_V6a 1000 E 33,809.00 33,808.95 0.00 33,808.95 0.00 0.03
LR1_DR02_VC02_V6a 1000 E 74,982.00 74,981.65 0.00 74,986.52 0.01 0.04
LR1_DR02_VC03_V7a 1000 E 40,446.00 40,340.01 -0.26 40,403.90 -0.10 0.05
LR1_DR02_VC03_V8a 1000 E 43,721.00 43,721.43 0.00 43,933.34 0.48 0.04
LR1_DR02_VC04_V8a 1000 E 41,657.00 41,708.69 0.12 41,847.57 0.46 0.08
LR1_DR02_VC05_V8a 1000 E 36,659.00 36,536.64 -0.33 36,616.09 -0.12 0.06
LR1_DR03_VC03_V10b 1000 M 92,941.00 92,810.27 -0.14 95,667.58 2.85 0.18
LR1_DR03_VC03_V13b 1000 E 124,921.00 125,094.57 0.14 127,203.29 1.79 0.43
LR1_DR03_VC03_V16a 1000 E 82,837.00 86,453.57 4.18 91,335.26 9.30 0.26
LR1_DR04_VC03_V15a 1000 E 73,312.00 74,072.61 1.03 74,714.08 1.88 0.67
LR1_DR04_VC03_V15b 1000 E 117,812.00 115,152.95 -2.31 116,451.06 -1.17 0.77
LR1_DR04_VC05_V17a 1000 E 72,876.00 73,446.65 0.78 73,864.03 1.34 0.89
LR1_DR04_VC05_V17b 1000 E 105,766.00 99,945.85 -5.82 103,534.03 -2.16 1.02
LR1_DR05_VC05_V25a 1000 E 105,328.00 105,579.09 0.24 105,995.54 0.63 2.57
LR1_DR05_VC05_V25b 1000 E 137,107.00 135,887.81 -0.90 140,504.31 2.42 2.87
LR1_DR08_VC05_V38a 1000 M 166,615.00 169,501.81 1.70 169,501.81 1.70 19.83
LR1_DR08_VC05_V40a 1000 E 178,593.00 185,214.23 3.57 185,214.23 3.57 24.94
LR1_DR08_VC05_V40b 100 M 200,746.00 222,644.62 9.84 234,070.81 14.24 4.46
LR1_DR08_VC10_V40a 100 M 185,538.00 188,966.37 1.81 189,991.43 2.34 3.43
LR1_DR08_VC10_V40b 100 M 206,315.00 220,751.77 6.54 234,282.23 11.94 4.40
LR1_DR12_VC05_V70a 100 H 278,647.00 287,636.40 3.13 289,736.67 3.83 18.37
LR1_DR12_VC05_V70b 100 M 308,555.00 316,041.88 2.37 322,301.84 4.27 19.98
LR1_DR12_VC10_V70a 10 H 283,154.00 289,629.21 2.24 290,673.25 2.59 12.61
LR1_DR12_VC10_V70b 10 H 295,126.00 300,647.25 1.84 304,041.16 2.93 15.71
Table 4: 180-period results
Best Best Average Average Time
Instance N Class Obj Total Cost Gap (%) Total Cost Gap (%) (h)
LR1_DR02_VC01_V6a 1000 E 52,167.00 52,167.21 0.00 52,167.21 0.00 0.07
LR1_DR02_VC02_V6a 1000 E 129,372.00 129,372.06 0.00 129,696.36 0.25 0.09
LR1_DR02_VC03_V7a 1000 M 60,547.00 61,302.02 1.23 61,950.19 2.27 0.11
LR1_DR02_VC03_V8a 1000 E 68,153.00 67,480.81 -1.00 70,478.74 3.30 0.09
LR1_DR02_VC04_V8a 1000 M 66,017.00 66,073.61 0.09 66,238.44 0.33 0.17
LR1_DR02_VC05_V8a 1000 M 58,619.00 58,087.60 -0.91 58,202.65 -0.72 0.14
LR1_DR03_VC03_V10b 1000 M 125,638.00 $1$4.95 122,230.31 -2.79 0.38
LR1_DR03_VC03_V13b 1000 E 165,764.00 159,667.18 -3.82 161,724.70 -2.50 0.90
LR1_DR03_VC03_V16a 1000 M 143,178.00 144,635.69 1.01 153,068.94 6.46 0.62
LR1_DR04_VC03_V15a 1000 M 118,621.00 118,360.01 -0.22 119,146.56 0.44 1.50
LR1_DR04_VC03_V15b 1000 E 189,989.00 187,064.87 -1.56 188,776.86 -0.64 1.66
LR1_DR04_VC05_V17a 1000 M 117,710.00 118,507.37 0.67 118,868.81 0.97 2.10
LR1_DR04_VC05_V17b 1000 M 159,168.00 152,523.12 -4.36 154,452.82 -3.05 2.27
LR1_DR05_VC05_V25a 1000 M 171,620.00 171,546.87 -0.04 172,104.96 0.28 6.04
LR1_DR05_VC05_V25b 1000 M 205,368.00 214,609.49 4.31 219,522.01 6.45 6.57
LR1_DR08_VC05_V38a 100 H 274,244.00 276,990.76 0.99 281,247.89 2.49 9.00
LR1_DR08_VC05_V40a 100 H 296,760.00 302,172.23 1.79 304,072.84 2.40 9.77
LR1_DR08_VC05_V40b 100 M 337,559.00 345,775.18 2.38 354,956.21 4.90 11.61
LR1_DR08_VC10_V40a 100 H 304,261.00 307,761.72 1.14 309,323.52 1.64 9.41
LR1_DR08_VC10_V40b 100 H 331,775.00 347,981.67 4.66 356,027.44 6.81 11.99
LR1_DR12_VC05_V70a 10 H 460,566.00 481,166.70 4.28 506,963.62 9.15 3.40
LR1_DR12_VC05_V70b 10 H 491,160.00 542,281.73 9.43 573,240.49 14.32 3.77
LR1_DR12_VC10_V70a 10 H 466,975.00 480,225.26 2.76 491,159.69 4.92 3.77
LR1_DR12_VC10_V70b 10 H 470,172.00 483,586.40 2.77 511,421.68 8.07 4.18
Table 5: 360-period results
Best Best Average Average Time
Instance N Class Obj Total Cost Gap (%) Total Cost Gap (%) (h)
LR1_DR02_VC01_V6a 1000 E 108,141.00 108,141.00 0.00 108,141.00 0.00 0.31
LR1_DR02_VC02_V6a 1000 E 283,031.00 281,693.77 -0.47 283,788.36 0.27 0.44
LR1_DR02_VC03_V7a 1000 H 124,282.00 124,772.19 0.39 126,940.04 2.09 0.50
LR1_DR02_VC03_V8a 1000 M 141,166.00 142,348.28 0.83 146,807.48 3.84 0.41
LR1_DR02_VC04_V8a 1000 M 138,693.00 138,888.24 0.14 139,237.84 0.39 0.84
LR1_DR02_VC05_V8a 1000 H 122,598.00 122,520.54 -0.06 122,859.71 0.21 0.71
LR1_DR03_VC03_V10b 1000 H 259,888.00 255,089.04 -1.88 260,807.85 0.35 1.56
LR1_DR03_VC03_V13b 1000 N 316,441.00 307,047.81 -3.06 312,621.28 -1.22 3.88
LR1_DR03_VC03_V16a 1000 N 327,793.00 338,115.45 3.05 349,216.16 6.13 3.56
LR1_DR04_VC03_V15a 1000 N 252,710.00 253,395.47 0.27 256,288.64 1.40 7.75
LR1_DR04_VC03_V15b 1000 N 339,308.00 336,787.96 -0.75 340,710.34 0,41 7.63
LR1_DR04_VC05_V17a 1000 H 251,623.00 253,554.51 0.76 254,040.64 0.95 11.23
LR1_DR04_VC05_V17b 1000 M 304,507.00 306,482.10 0.64 312,046.69 2.42 10.18
LR1_DR05_VC05_V25a 100 H 368,628.00 372,314.12 0.99 373,871.33 1.40 8.84
LR1_DR05_VC05_V25b 100 H 410,053.00 472,870.66 13.28 494,627.59 17.10 9.79
LR1_DR08_VC05_V38a 100 H 596,969.00 683,046.79 12.60 683,046.79 12.60 22.52
LR1_DR08_VC05_V40a 100 H 652,380.00 654,769.74 0.36 659,967.40 1.15 20.30
LR1_DR08_VC05_V40b 100 H 709,713.00 779,769.69 8.98 804,362.42 11.77 22.41
LR1_DR08_VC10_V40a 100 H 663,245.00 670,198.99 1.04 672,938.28 1.44 19.92
LR1_DR08_VC10_V40b 100 H 724,513.00 786,302.74 7.86 794,420.68 8.80 21.26
LR1_DR12_VC05_V70a 10 H 1,021,389.00 1,125,724.51 9.27 1,193,627.62 14.43 20.07
LR1_DR12_VC05_V70b 10 H 1,093,013.00 1,378,324.17 20.70 1,435,397.61 23.85 21.35
LR1_DR12_VC10_V70a 10 H 1,024,399.00 1,043,792.70 1.86 1,046,678.91 2.13 19.39
LR1_DR12_VC10_V70b 10 H 1,001,541.00 1,229,992.15 18.57 1,306,455.89 23.34 16.12
@@ 5.3.2 BS and Greedy Heuristic Results
As mentioned in Section 4.4, different combinations of nondeterministic approaches were tested for the greedy heuristic: applying it separately to port selection, vessel selection, both or none. For these tests, the 13 smallest Group 2 instances were used, with four different seeds, all with a 120-day horizon, and $q$ was set to 5. The size of the instances allowed to set a large value for $N$, 1000. The configuration that achieved the best result, with the lowest aggregate average value among instances, was to apply randomness only to port selection while keeping the selection of the vessel deterministic. The second best was to apply randomness in both decisions followed by keeping greedy heuristic deterministic. The application of randomness only on vessel selection was the last. Table 9 (A) presents the complete results of these tests.
Once it was established that nondeterministic methods would be used solely for the port, additional tests were performed to evaluate three distinct techniques for aggregating the results of the greedy algorithm: minimum, mean, or median cost. The experiments were then conducted for the 13 smallest instances using three horizon sizes. For each instance, four executions with different seeds were carried out. The number of nodes allowed per level in the BS was set to 200. This decrease is attributed to the fact that instances now take into account a time horizon extending up to 360 days. The method that showed the best performance was the median with the lowest aggregate average value among the instances. The second best was the mean followed by the minimum. The complete results of these tests are shown in Table 10 (A).
Table 6 highlights the importance of the BS within the algorithm. Each column corresponds to a specific algorithm stage (After BS, After LS, and After ILS), allowing a comparison of the impact of BS on the results. The percentages indicate the improvement achieved by incorporating BS as opposed to relying solely on the greedy heuristic. Comparing the results obtained through the exclusive use of greedy heuristic versus the application of the BS algorithm, it is possible to observe that the BS approach yields an average improvement of 18.40% in the solution following the BS stage. In particular, local search has a significant impact on results in the absence of BS, reducing the advantage of employing BS to 7.14%. When the complete heuristic is applied, the benefit of BS decreases further to 3.57% for this set of instances. However, its use remains beneficial, despite the increased computational effort.